# arXiv:2603.00425v1[cs.LG] 28 Feb 2026

### 28 Feb 2026 v1

@@ 1 Introduction
Despite their empirical success, activation steering methods remain largely heuristic-driven. Current research in this area primarily relies on exhaustive trial-and-error to determine critical design choices, such as optimal parameterization and intervention sites. For instance, JoLA lai2025joint treats the intervention site as a learnable parameter and offers little insight into the underlying mechanics. Furthermore, most prominent methods evaluate specific intervention points in isolation, lacking a comparative framework to explain why certain loci outperform others. Consequently, the design of steering interventions remains a black box process lacking generalizable principles.
@@ 2 Related Work
Parameter Efficient Fine-tuning Standard PEFT methods reduce adaptation costs by training small weight-space modules. This includes both approaches like classic bottleneck adapters houlsby2019parameter, which insert trainable layers into the transformer block, and reparameterization approaches like LoRA hu2022lora and its variants (DoRA liu2024dora, QLoRA dettmers2023qlora). While they significantly reduce fine-tuning costs, such weight-space modifications do not seek to use intermediate activations—the key component in steering.
Activation Steering. Activation steering intervenes directly on intermediate activations during the forward pass, often achieving parameter savings an order of magnitude greater than weight-based adapters. Initial research focused on identifying latent subspaces for specific traits, such as truthfulness li2023inference, style rimsky2024steering, function execution todd2023function, or alignment adila2024free, and steer the model towards the desired subspaces. More recently, methods like ReFT wu2024reft, LoFiT yin2024lofit, and JoLA lai2025joint have transitioned toward trainable interventions that minimize a standard fine-tuning loss.
However, these methods remain largely heuristic-driven; design choices regarding intervention locus and parameterization are typically determined through empirical search rather than mechanical understanding. While JoLA lai2025joint attempts to unify these heuristics by treating design choices as hyperparameters to be optimized, this increases optimization complexity and does not change the black-box nature of steering design.
Interplay between adaptation paradigms. Recent research has sought to unify different adaptation paradigms through a mechanistic lens. Significant effort has been directed toward interpreting in-context learning (ICL) as a form of implicit weight fine-tuning dai-etal-2023-gpt, von2023transformers, zhang2024trained, with recent evidence suggesting these implicit updates are effectively rank-1 dherin2025learning. This provides a direct bridge to activation steering, as an activation update can be viewed as a rank-1 LoRA update constrained to a single direction. Furthering this connection, bigelow2025belief established a link between ICL and steering activations toward pre-identified behavioral subspaces.
@@ 3.1 Setup and Notation
We use $\Delta\cdot$ for a small change to a variable, while $\delta\cdot$ is a learned update to a model parameter. Steering updates a model’s activations by updating $h$ to $h\to h+\delta h$. The update $\delta h$, called an adapter, can depend on the input $h$. For example, in ReFT wu2024reft, $\delta h$ is a specific, low-rank linear function acting on $h$. An alternative is a low-rank autoencoder, $\delta h=W_{2}\phi(W_{1}h)$, where $W_{2}\in\mathbb{R}^{d\times r}$ and $W_{1}\in\mathbb{R}^{r\times d}$. Fine-tuning is an update to the weights of a model, replacing $W$ with $W+\delta W$. The parameter $\delta W$ is a learned constant.
@@ 3.2 First Order Analysis
First, we study the relationship between steering and fine-tuning in order to extract insights into where steering interventions should take place. Consider the Gated Linear Unit (GLU), a common variant of MLP dauphin2017language, shazeer2020glu used in most modern LLMs, such as Llama, Gemma, and Qwen. This module is parameterized as $\mathrm{GLU}(h)=W_{d}(\phi(W_{g}h)\odot W_{u}h).$
@@ 3.3 Post-Block and Post-MLP
With the similarity between steering and fine-tuning understood, we investigate how closely a linear post-block steering update can mimic a ReFT-style post-MLP steering.
@@
This proposition tells us that the ability for a post-block steering to match a post-MLP steering depends on how similar the output of the GLU and the output of $h+\mathrm{Attn}(h)$ are. The greater the difference, the greater the principal angles Stewart90 will be, and therefore the greater the error. When the skip connection is relatively small, these two subspaces will be unlikely to differ much.
@@ 4.1 Joint Learning in Weight and Activation Space
The relationship between the left-singular vectors of $RY^{\top}$ and $RF^{\top}$ is not guaranteed. In fact, since $F$ is an arbitrary function, these two can have incredibly dissimilar dominant singular vector spaces. However, for real data such as Winogrande sakaguchi2021winogrande, Figure 3 shows the inner products for the top 50 singular vectors of these matrices at three layers of the Llama model being trained.
@@ 5.1 Post-Block Steering
where $W_{1}\in\mathbb{R}^{r\times d}$ and $W_{2}\in\mathbb{R}^{d\times r}$ project the hidden state $h$ into a low-dimensional subspace of rank $r$ and project it back to the model’s dimension. We evaluate both linear and non-linear variants, where $\phi$ is either the identity function or a non-linear activation (e.g., SiLU). While structurally similar to ReFT wu2024reft, we introduce two principled improvements derived from Section 3:
  • Theoretically-justified Locus: We shift the intervention site to the post-block output (Figure 1, right), a choice dictated by our derivation of the weight-activation map.
@@ 6 Experiments
  • Principled steering approximates full-parameter fine-tuning (Section 6.1). Post-block steering performs within a $0.4\%\text{--}1.5\%$ margin of full-parameter fine-tuning (SFT) while training only $0.04\%$ model parameters.
  • Generalization to complex training paradigms (Section 6.2). The effectiveness of our approach extends to more complex optimization landscapes like instruction tuning.
@@
We evaluate our approach on six benchmarks: three commonsense reasoning datasets, BoolQ clark2019boolq, Winogrande sakaguchi2021winogrande, ARC Challenge allenai:arc; two mathematical reasoning, GSM8K cobbe2021training and AQuA ling-etal-2017-program; and one long-context task, ListOps from Long Range Arena tay2020long, nangia2018listops. Prompt templates and dataset details is available in Appendix F.1.
@@ 6.1 Approximating full-parameter SFT
    Performance-oriented PEFT: We compare against LoRA hu2022lora, the industry standard for weight-space adaptation, and ReFT wu2024reft, a high-performance representation steering method (Table 1).
@@
    Ultra-Efficient Steering: We benchmark against LoFiT yin2024lofit and JoLA lai2025joint, which target extreme parameter efficiency (Table 2).
The parameter budget (Params $(\%$)) is measured as the percentage of trainable parameters relative to the total model size. To ensure a fair comparison, we match the parameter counts of most baselines by adjusting our adapter rank. We allocate LoRA a larger budget ($0.45\%$ vs ours $0.04\%$) because its performance degrades significantly at lower ranks. For the variants shown in Table 2, Ours (1-vec) denotes a version of our method that trains single steering vectors instead of matrices. All steering adapters in this subsection are linear ($\phi$ is the identity function). For a fair comparison, ReFT is also applied at all layers (matching our global intervention policy), so that the only difference between the two methods is the intervention locus (post-MLP vs. post-block).
@@ Results.
Table 1 shows that our approach maintains an average performance gap within $0.4\%$$1.5\%$ of SFT while training only $0.04\%$ parameters. On 1B-scale models, we match or outperform LoRA despite training $11\times$ fewer parameters. Compared to ReFT at the same $0.04\%$ budget, our method is significantly more stable on complex tasks. The performance gap is particularly pronounced on ListOps, a long-range dependency task where attention plays a critical role. While ReFT drops by up to $16.9\%$, our approach limits the gap to $3.1\%$. This aligns with our theoretical analysis (Takeaway 2): post-block steering captures attention effects that post-MLP methods cannot access.
@@ 6.2 Generalization to more complex objectives
To test the generalization of post-block steering to complex, open-ended generation, we perform instruction tuning using Llama 3.1 8B. We train on Alpaca-Cleaned (alpaca) (${\sim}52$K pairs) for 3 epochs and compare against full-parameter SFT, LoRA ($r=8$), and ReFT. For ReFT, we evaluate both its published configuration ($r=4$ at layers 3, 9, 18, and 24) and a parameter-matched variant ($r=16$ at all layers). While we report the best of two learning rates for SFT and LoRA, our method uses a fixed learning rate of $2\times 10^{-3}$. All methods use an effective batch size of $16$. Models are evaluated via AlpacaEval 2.0 (alpaca_eval), reporting length-controlled (LC) win rates against GPT-4 Turbo and standard errors (SE) across 805 instructions.
@@ Results.
Table 3 summarizes the instruction tuning performance. Post-block steering with a nonlinear adapter achieves an LC win rate of 11.34%, falling within 0.15% gap of full-parameter SFT (11.49%). Notably, both our linear variant (11.00%) and the parameter-matched ReFT baseline (10.96%) exhibit performance comparable to SFT, with both our adapters falling within one standard error of the SFT mean. In contrast, LoRA achieves a lower win rate (10.52%) despite a significantly larger parameter budget ($0.26\%$). These results confirm that the extreme parameter efficiency of post-block steering generalizes to the challenging regime of open-ended instruction following. Furthermore, the marginal 0.34% gap between our linear and nonlinear variants suggests that while nonlinear parameterization offers peak performance, the linear approximation is highly robust, a trend we examine more broadly next.
Table 4: Joint training vs. its components. “Joint" is naive joint training, “Joint-Orth" incorporates the orthogonality constraint. Joint uses LoRA (0.45%) + Adapter (0.34%), total 0.79% trainable params. LoRA $r=16$ has $0.9\%$ trainable params.
@@ 6.3 Effect of nonlinearity
A natural concern is whether our first-order analysis holds under practical training conditions. To test this, we examine how non-linearity interacts with increasing rank and model scale using Qwen3 yang2025qwen3 model suite (0.6B, 1.7B, 4B, and 8B) on Winogrande.
@@ C.1 The Specific Location
Other methods, such as ReFT, steer the outputs of the MLP or the attention modules before the skip-connections are added. In notation, for a steering vector $\delta h$,
@@
In principle, these can be identical for a completely freely parameterized $\delta h$. However, there is an important difference which is that $\delta h$ can only depend on parts of the model and not others. The ReFT location has $\delta h(\mathrm{GLU}(h+\mathrm{Attn}(h)))$ while our method has $\delta h(y(h))$. To investigate this further in a clean way, take both steering vectors as a linear update from the existing hidden state, i.e.
$\delta h_{\mathrm{ReFT}}=A_{\mathrm{ReFT}}h$
@@
Note that these do not necessarily have to coincide, based on the the down projection $W_{d}$ in the $\mathrm{GLU}$. In fact, assume that $A\cap B=\{0\}$, and denote the projections onto these subspaces, remove all components from the other as $P_{A}$ and $P_{B}$. Then, for any $A_{\mathrm{ReFT}}$,
$A_{\mathrm{steer}}=A_{\mathrm{ReFT}}P_{B}$
@@
$y_{\mathrm{ReFT}}(h)=y_{\mathrm{steer}}(h)$
@@
$=h+\mathrm{Attn}(h)+(\mathrm{GLU}(h+\mathrm{Attn}(h))+A_{\mathrm{ReFT}}P_{B}(h+\mathrm{Attn}(h))+A_{\mathrm{ReFT}}P_{B}(\mathrm{GLU}(h+\mathrm{Attn}(h)))$
$=h+\mathrm{Attn}(h)+(\mathrm{GLU}(h+\mathrm{Attn}(h))+A_{\mathrm{ReFT}}(\mathrm{GLU}(h+\mathrm{Attn}(h)))$
$=h+\mathrm{Attn}(h)+(\mathrm{GLU}(h+\mathrm{Attn}(h)+\delta h_{\mathrm{ReFT}})$
@@
Let $V$ and $V^{\prime}$ be the right-singular matrices of $h+\mathrm{Attn}(h)+\mathrm{GLU}(h+\mathrm{Attn}(h))$ and $A_{\mathrm{ReFT}}\mathrm{GLU}(h+\mathrm{Attn}(h))$ respectively, and let $X,Y$ be such that $Y_{i}=h_{i}+\mathrm{Attn}(h_{i})$ and $X_{i}=\mathrm{GLU}(Y_{i})$. Then, the optimal relative error satisfies
$\min_{A}\frac{\|A(X+Y)-A_{\mathrm{ReFT}}X\|_{F}^{2}}{\|A_{\mathrm{ReFT}}X\|_{F}^{2}}=\sum_{i=1}^{d}\left(\frac{(\sigma^{\prime})_{i}^{2}}{\sum_{j=1}^{d}(\sigma^{\prime})_{j}^{2}}\right)\sin^{2}\theta_{i}$
where $\sigma_{i}$ is the $i$-th singular value of $A_{\mathrm{ReFT}}X$ and $\theta_{i}$ is the $i$-th principle angle between $V$ and $V^{\prime}$.
@@
For some $A_{\mathrm{ReFT}}$, the best possible steering matrix $A_{\mathrm{steer}}$ will be
$A_{\mathrm{steer}}=\operatorname*{argmin}_{A}\mathcal{L}(A)=\operatorname*{argmin}_{A}\frac{\|A(X+Y)-A_{\mathrm{ReFT}}X\|_{F}^{2}}{\|A_{\mathrm{ReFT}}X\|_{F}^{2}}$
@@
$\mathcal{L}(A_{\mathrm{steer}})=\frac{1}{n}\|A_{\mathrm{ReFT}}X((X+Y)^{\top}((X+Y)(X+Y)^{\top})^{-1}(X+Y)-I)\|_{F}^{2}$
@@
$\mathcal{L}(A_{\mathrm{steer}})=\frac{\|A_{\mathrm{ReFT}}X(I-VV^{\top})\|_{F}^{2}}{\|A_{\mathrm{ReFT}}X\|_{F}^{2}}$
@@
$U^{\prime}\Sigma^{\prime}(V^{\prime})^{\top}=A_{\mathrm{ReFT}}X$
be the SVD of $A_{\mathrm{ReFT}}X$, we can decompose the numerator of $\mathcal{L}$ as
$\|A_{\mathrm{ReFT}}X(I-VV^{\top})\|_{F}^{2}$ $=\sum_{i=1}^{d}(\sigma^{\prime})_{i}^{2}\|(I-VV^{\top})V^{\prime}_{i}\|_{2}^{2}$

### 06 Mar 2026 v2

# Weight Updates as Activation Shifts: A Principled Framework for Steering

@@ 1 Introduction
Despite their empirical success, activation steering methods remain largely heuristic-driven. Current research in this area primarily relies on exhaustive trial-and-error to determine critical design choices, such as optimal parameterization and intervention sites. For instance, JoLA (lai2025joint) treats the intervention site as a learnable parameter and offers little insight into the underlying mechanics. Furthermore, most prominent methods evaluate specific intervention points in isolation, lacking a comparative framework to explain why certain loci outperform others. Consequently, the design of steering interventions remains a black box process lacking generalizable principles.
@@ 2 Related Work
Parameter Efficient Fine-tuning Standard PEFT methods reduce adaptation costs by training small weight-space modules. This includes both approaches like classic bottleneck adapters (houlsby2019parameter), which insert trainable layers into the transformer block, and reparameterization approaches like LoRA (hu2022lora) and its variants (DoRA (liu2024dora), QLoRA (dettmers2023qlora)). While they significantly reduce fine-tuning costs, such weight-space modifications do not seek to use intermediate activations—the key component in steering.
Activation Steering. Activation steering intervenes directly on intermediate activations during the forward pass, often achieving parameter savings an order of magnitude greater than weight-based adapters. Initial research focused on identifying latent subspaces for specific traits, such as truthfulness (li2023inference), style (rimsky2024steering), function execution (todd2023function), or alignment (adila2024free), and steer the model towards the desired subspaces. More recently, methods like ReFT (wu2024reft), LoFiT (yin2024lofit), and JoLA (lai2025joint) have transitioned toward trainable interventions that minimize a standard fine-tuning loss.
However, these methods remain largely heuristic-driven; design choices regarding intervention locus and parameterization are typically determined through empirical search rather than mechanical understanding. While JoLA (lai2025joint) attempts to unify these heuristics by treating design choices as hyperparameters to be optimized, this increases optimization complexity and does not change the black-box nature of steering design.
Interplay between adaptation paradigms. Recent research has sought to unify different adaptation paradigms through a mechanistic lens. Significant effort has been directed toward interpreting in-context learning (ICL) as a form of implicit weight fine-tuning (dai-etal-2023-gpt, von2023transformers, zhang2024trained), with recent evidence suggesting these implicit updates are effectively rank-1 (dherin2025learning). This provides a direct bridge to activation steering, as an activation update can be viewed as a rank-1 LoRA update constrained to a single direction. Furthering this connection, (bigelow2025belief) established a link between ICL and steering activations toward pre-identified behavioral subspaces.
@@ 3.1 Setup and Notation
We use $\Delta\cdot$ for a small change to a variable, while $\delta\cdot$ is a learned update to a model parameter. Steering updates a model’s activations by updating $h$ to $h\to h+\delta h$. The update $\delta h$, called an adapter, can depend on the input $h$. For example, in ReFT (wu2024reft), $\delta h$ is a specific, low-rank linear function acting on $h$. An alternative is a low-rank autoencoder, $\delta h=W_{2}\phi(W_{1}h)$, where $W_{2}\in\mathbb{R}^{d\times r}$ and $W_{1}\in\mathbb{R}^{r\times d}$. Fine-tuning is an update to the weights of a model, replacing $W$ with $W+\delta W$. The parameter $\delta W$ is a learned constant.
@@ 3.2 First Order Analysis
First, we study the relationship between steering and fine-tuning in order to extract insights into where steering interventions should take place. Consider the Gated Linear Unit (GLU), a common variant of MLP (dauphin2017language, shazeer2020glu) used in most modern LLMs, such as Llama, Gemma, and Qwen. This module is parameterized as $\mathrm{GLU}(h)=W_{d}(\phi(W_{g}h)\odot W_{u}h).$
@@ 3.3 Post-Block and Post-MLP
With the similarity between steering and fine-tuning understood, we investigate how closely a linear post-block steering update can mimic post-MLP steering.
@@
This proposition tells us that the ability for a post-block steering to match a post-MLP steering depends on how similar the output of the GLU and the output of $h+\mathrm{Attn}(h)$ are. The greater the difference, the greater the principal angles (Stewart90) will be, and therefore the greater the error. When the skip connection is relatively small, these two subspaces will be unlikely to differ much.
@@ 4.1 Joint Learning in Weight and Activation Space
The relationship between the left-singular vectors of $RY^{\top}$ and $RF^{\top}$ is not guaranteed. In fact, since $F$ is an arbitrary function, these two can have incredibly dissimilar dominant singular vector spaces. However, for real data such as Winogrande (sakaguchi2021winogrande), Figure 3 shows the inner products for the top 50 singular vectors of these matrices at three layers of the Llama model being trained.
@@ 5.1 Post-Block Steering
where $W_{1}\in\mathbb{R}^{r\times d}$ and $W_{2}\in\mathbb{R}^{d\times r}$ project the hidden state $h$ into a low-dimensional subspace of rank $r$ and project it back to the model’s dimension. We evaluate both linear and non-linear variants, where $\phi$ is either the identity function or a non-linear activation (e.g., SiLU). While structurally similar to ReFT (wu2024reft), we introduce three principled improvements derived from Section 3:
  • Theoretical Foundation of the Locus: While ReFT identifies the residual stream as a high-performance intervention site through empirical selection, our framework provides a formal theoretical foundation for this choice (Section 3.2).
@@
  • Bottleneck Adapter Parameterization: ReFT focuses on a specific low-rank linear parameterizations with constraints such as orthonormal rows. We employ a more flexible bottleneck adapter without these constraints and permits non-linear transformations.
@@ 5.2 Orthogonality-Constrained Joint Adaptation
Table 4: RL on DeepSeek-R1-Distill-Qwen-1.5B with GSM8K.
Method Params (%) Pass@1
Base model - 10.2
LoRA 0.52% $81.5\pm 0.7$
Ours (nonlinear) 0.04% $84.7\pm 0.5$
Ours (linear) 0.04% $84.3\pm 0.8$
@@ 6 Experiments
  • Principled steering approximates full-parameter fine-tuning (Section 6.1). Post-block steering performs within a $0.2\%\text{--}0.9\%$ margin of full-parameter fine-tuning (SFT) while training only $0.04\%$ model parameters.
  • Generalization to complex training paradigms (Section 6.2). The effectiveness of our approach extends to more complex optimization landscapes like instruction tuning and reinforcement learning (RL).
@@
We evaluate our approach on six benchmarks: three commonsense reasoning datasets, BoolQ (clark2019boolq), Winogrande (sakaguchi2021winogrande), ARC Challenge (allenai:arc); two mathematical reasoning, GSM8K (cobbe2021training) and AQuA (ling-etal-2017-program); and one long-context task, ListOps from Long Range Arena (tay2020long, nangia2018listops). Prompt templates and dataset details is available in Appendix F.1.
@@ 6.1 Approximating full-parameter SFT
    Performance-oriented PEFT: We compare against LoRA (hu2022lora), the industry standard for weight-space adaptation, and ReFT (wu2024reft), a high-performance representation steering method (Table 1).
@@
    Ultra-Efficient Steering: We benchmark against LoFiT (yin2024lofit) and JoLA (lai2025joint), which target extreme parameter efficiency (Table 2).
The parameter budget (Params $(\%$)) is measured as the percentage of trainable parameters relative to the total model size. To ensure a fair comparison, we match the parameter counts of most baselines by adjusting our adapter rank. We allocate LoRA a larger budget ($0.45\%$ vs ours $0.04\%$) because its performance degrades significantly at lower ranks. For the variants shown in Table 2, Ours (1-vec) denotes a version of our method that trains single steering vectors instead of matrices. All steering adapters in this subsection are linear ($\phi$ is the identity function). For a fair comparison, ReFT is also applied at all layers (matching our global intervention policy).
@@ Results.
Table 1 shows that our approach maintains an average performance gap within $0.2\%$$0.9\%$ of SFT while training only $0.04\%$ parameters. On 1B-scale models, we match or outperform LoRA despite training $11\times$ fewer parameters. Compared to ReFT at the same $0.04\%$ budget, our method is significantly more stable on complex tasks. The performance gap is particularly pronounced on ListOps, a long-range dependency task where attention plays a critical role. While ReFT drops by up to $16.9\%$, our approach limits the gap to $3.1\%$.
@@ 6.2 Generalization to more complex objectives
Next, we evaluate post-block steering on instruction tuning and RL, testing whether our results generalizes from the structured reasoning tasks studied previously to open-ended generation and iterative policy optimization.
For instruction tuning, we fine-tune Llama 3.1 8B on Alpaca-Cleaned (alpaca) (${\sim}52$K pairs) for 3 epochs. We compare post-block steering against full-parameter SFT, LoRA ($r=8$ and $r=16$), and two ReFT variants: the published sparse configuration ($r=4$ at 4 layers) and a parameter-matched variant ($r=16$ at all layers). While we report the best of two learning rates for SFT and LoRA, our method uses a fixed learning rate of $2\times 10^{-3}$. Models are evaluated on AlpacaEval 2.0 (alpaca_eval) reporting length-controlled (LC) win rates against GPT-4 Turbo and standard errors (SE) across 805 instructions.
For RL, we perform Group Relative Policy Optimization (GRPO) (shao2024deepseekmath) on DeepSeek-R1-Distill-Qwen-1.5B (deepseekai2025deepseekr1incentivizingreasoningcapability), comparing our post-block adapter against the standard LoRA implementation in TRL (vonwerra2020trl, schulman2025lora). Both methods use rank $r=8$.
@@ Results.
Table 3 summarizes the instruction tuning results. Our nonlinear post-block adapter achieves an LC win rate of 11.34%, nearly closing the gap to full-parameter SFT (11.49%) within 0.15%. Both our linear (11.00%) and nonlinear variants outperform LoRA (10.52%), despite LoRA using more trainable parameters. This confirms that the expressivity of post-block steering extends to open-ended instruction following.
[…]
We also note that the marginal gap between our linear and nonlinear variants on these complex scenarios suggests that while nonlinear parameterization offers peak performance, the linear approximation is highly robust, a trend we examine more broadly next in Section 6.3.
Table 5: Joint training vs. its components. “Joint” is naive joint training, “Joint-Orth” incorporates the orthogonality constraint. Joint uses LoRA (0.45%) + Adapter (0.34%), total 0.79% trainable params. LoRA $r=16$ has $0.9\%$ trainable params.
@@ 6.3 Effect of nonlinearity
A natural concern is whether our first-order analysis holds under practical training conditions. To test this, we examine how non-linearity interacts with increasing rank and model scale using Qwen3 (yang2025qwen3) model suite (0.6B, 1.7B, 4B, and 8B) on Winogrande.
@@ C.1 The Specific Location
Other methods steer the outputs of the MLP or the attention modules before the skip-connections are added. In notation, for a steering vector $\delta h$,
@@
In principle, these can be identical for a completely freely parameterized $\delta h$. However, there is an important difference which is that $\delta h$ can only depend on parts of the model and not others. Post MLP methods location has $\delta h(\mathrm{GLU}(h+\mathrm{Attn}(h)))$ while our method has $\delta h(y(h))$. To investigate this further in a clean way, take both steering vectors as a linear update from the existing hidden state, i.e.
$\delta h_{p}=A_{p}h$
@@
Note that these do not necessarily have to coincide, based on the the down projection $W_{d}$ in the $\mathrm{GLU}$. In fact, assume that $A\cap B=\{0\}$, and denote the projections onto these subspaces, remove all components from the other as $P_{A}$ and $P_{B}$. Then, for any $A_{p}$,
$A_{\mathrm{steer}}=A_{p}P_{B}$
@@
$y_{p}(h)=y_{\mathrm{steer}}(h)$
@@
$=h+\mathrm{Attn}(h)+(\mathrm{GLU}(h+\mathrm{Attn}(h))+A_{p}P_{B}(h+\mathrm{Attn}(h))+A_{p}P_{B}(\mathrm{GLU}(h+\mathrm{Attn}(h)))$
$=h+\mathrm{Attn}(h)+(\mathrm{GLU}(h+\mathrm{Attn}(h))+A_{p}(\mathrm{GLU}(h+\mathrm{Attn}(h)))$
$=h+\mathrm{Attn}(h)+(\mathrm{GLU}(h+\mathrm{Attn}(h)+\delta h_{p})$
@@
$\min_{A}\frac{\|A(X+Y)-A_{p}X\|_{F}^{2}}{\|A_{p}X\|_{F}^{2}}=\sum_{i=1}^{d}\left(\frac{(\sigma^{\prime})_{i}^{2}}{\sum_{j=1}^{d}(\sigma^{\prime})_{j}^{2}}\right)\sin^{2}\theta_{i}$
where $\sigma_{i}$ is the $i$-th singular value of $A_{p}X$ and $\theta_{i}$ is the $i$-th principle angle between $V$ and $V^{\prime}$.
@@
For some $A_{p}$, the best possible steering matrix $A_{\mathrm{steer}}$ will be
$A_{\mathrm{steer}}=\operatorname*{argmin}_{A}\mathcal{L}(A)=\operatorname*{argmin}_{A}\frac{\|A(X+Y)-A_{p}X\|_{F}^{2}}{\|A_{p}X\|_{F}^{2}}$
@@
$\mathcal{L}(A_{\mathrm{steer}})=\frac{1}{n}\|A_{p}X((X+Y)^{\top}((X+Y)(X+Y)^{\top})^{-1}(X+Y)-I)\|_{F}^{2}$
@@
$\mathcal{L}(A_{\mathrm{steer}})=\frac{\|A_{p}X(I-VV^{\top})\|_{F}^{2}}{\|A_{p}X\|_{F}^{2}}$
@@
$U^{\prime}\Sigma^{\prime}(V^{\prime})^{\top}=A_{p}X$
be the SVD of $A_{p}X$, we can decompose the numerator of $\mathcal{L}$ as
$\|A_{p}X(I-VV^{\top})\|_{F}^{2}$ $=\sum_{i=1}^{d}(\sigma^{\prime})_{i}^{2}\|(I-VV^{\top})V^{\prime}_{i}\|_{2}^{2}$